# RouteLLM: Learning to Route LLMs with Preference Data

### v1

@@ 4.1 Results
$\mathcal{D}_{\text{arena}}$ augmented BERT 19.58% 34.02% 0.751 (+50.2%)
with $\mathcal{D}_{\text{judge}}$ Causal LLM 31.50% 48.75% 0.679 (+35.8%)
@@
$\mathcal{D}_{\text{arena}}$ augmented BERT 41.30% 72.20% 0.572 (+14.4%)
with $\mathcal{D}_{\text{gold}}$ Causal LLM 35.49% 70.31% 0.600 (+19.9%)
@@
$\mathcal{D}_{\text{arena}}$ augmented BERT 44.76% 79.09% 0.531 (+6.9%)
with $\mathcal{D}_{\text{judge}}$ Causal LLM 33.64% 63.26% 0.622 (+25.3%)
@@ 4.3 Generalizing to other model pairs
$\mathcal{D}_{\text{arena}}$ augmented BERT 36.28% 50.83% 0.618 (+25.2%)
with $\mathcal{D}_{\text{judge}}$ Causal LLM 40.46% 55.83% 0.625 (+26.5%)
@@ Appendix E Independent Benchmarks
In Figure 2, we present the performance of our best-performing routers on MT Bench as compared to Unify AI [3] and notdiamond-001 [13], two other LLM routing systems released by companies. Here, we route between gpt-4-turbo-2024-04-09 [20] and Mixtral 8x7B [17] instead. For Unify AI, we select the best-performing router configuration on the user dashboard and use it for benchmarking. For notdiamond-001, we replace calls to GPT 3.5 with calls to Mixtral 8x7B. Both the matrix factorization router and causal LLM routers perform competitively when trained on $\mathcal{D}_{\text{arena}}$ augmented with $\mathcal{D}_{\text{judge}}$, achieving much stronger results that these existing routers for the same number of calls routed to GPT-4.

### v2

@@ 4.1 Results
$\mathcal{D}_{\text{arena}}+\mathcal{D}_{\text{judge}}$ BERT 19.58% 34.02% 0.751 (+50.2%)
Causal LLM 31.50% 48.75% 0.679 (+35.8%)
@@
$\mathcal{D}_{\text{arena}}+\mathcal{D}_{\text{gold}}$ BERT 41.30% 72.20% 0.572 (+14.4%)
Causal LLM 35.49% 70.31% 0.600 (+19.9%)
@@
$\mathcal{D}_{\text{arena}}+\mathcal{D}_{\text{judge}}$ BERT 44.76% 79.09% 0.531 (+6.9%)
Causal LLM 33.64% 63.26% 0.622 (+25.3%)
@@ 4.3 Generalizing to other model pairs
$\mathcal{D}_{\text{arena}}+\mathcal{D}_{\text{judge}}$ BERT 36.28% 50.83% 0.618 (+25.2%)
Causal LLM 40.46% 55.83% 0.625 (+26.5%)
@@ Appendix E Independent Benchmarks
In Figure 2, we present the performance of our best-performing routers on MT Bench as compared to Unify AI [3] and Martian [19], two existing commercial offerings for LLM routing. Here, we route between gpt-4-turbo-2024-04-09 [20] as our strong model, and mixtral-8x7b-instruct-v0.1 [16] or llama-2-70b-chat [26] as our weak model depending on the available models. For Unify AI, we select the best-performing router configuration on the user dashboard and use it for benchmarking. For Martian, we optimize for performance and specify the maximum cost per million tokens as \$10.45, approximating this value using public inference costs [21, 24] based on a 1:1 input:output token ratio such that 50% of calls are routed to GPT-4. Both the matrix factorization router and causal LLM routers perform very competitively when trained on $\mathcal{D}_{\text{arena}}$ augmented with $\mathcal{D}_{\text{judge}}$, achieving the same performance as these routers with up to 40% fewer calls routed to GPT-4.